# Dynamics of quantum Fisher and Wigner-Yanase skew information following a noisy quench

### v1

@@ Abstract
We study the influence of noise on the dynamics of a transverse field Ising model when quenched across a quantum critical point. To quantify two‐spin correlations properties, we employ the quantum Fisher information (QFI) and Wigner-Yanase skew information (WYSI) as measures of quantum coherence. In the absence of noise, despite the entanglement, both QFI and WYSI increase monotonically with the ramp quench time, approaching their adiabatic limits without exhibiting any Kibble–Zurek–type scaling with quench duration. When noise is added to the quench protocol, the coherence dynamics change dramatically: QFI and WYSI both decay exponentially with time scale of a ramp quench, with the exponent depending on the noise intensity. Furthermore, the maximum ramp time, at which either of these measures reach their maximum, scales linearly with the noise variance, featuring the same exponent that determines the optimal annealing time for minimizing defect production in noisy quantum annealing.
@@ I Introduction
Quantum phase transitions (QPTs) have been one of the most important topics in the area of strongly correlated systems [1, 2, 3]. There have been recent developments in theoretical investigation of QPTs from the viewpoint of quantum information theory. Quantities like concurrence [4, 5], entanglement entropy [6, 7], fidelity susceptibility [8, 9, 10, 11] and geometric phases [12, 13] have been shown to capture the non-analyticities associated with a quantum critical point (QCP). Recent advances in the studies of ultracold atoms trapped in optical lattices have opened a new avenue of investigation of the non-equilibrium dynamics of QPTs in quantum many body systems [14, 15]. Divergences in correlation length and time scales at QCPs introduce non-trivial effects in the dynamics of a quantum system near a QCP. In particular, if a parameter of the Hamiltonian is tuned slowly so as to drive the system through a QCP, the dynamics fails to be adiabatic for any finite rate of variation of the parameter. This non-adiabaticity in the response of the system eventually leads to generation of defects in the final state. A quantitative analysis of defect production in a system following a quench can be carried out using the Kibble-Zurek (KZ) theory [16, 17] extended to quantum spin chains [18, 19, 20, 21, 22, 23].
@@
This raises the fundamental questions about the generation of quantum correlations following a noisy quench. Have the quantum correlations generated under noisy quench via QCPTs? If so, what is the effect from noise on the dynamical behavior? To address this question we invoke the dynamics of quantum Fisher information (QFI) and Wigner-Yanase skew information (WYSI) in the noiseless and noisy driven transverse field Ising model (TFI). QFI [53] is an extension of Fisher information [54] to the quantum realm, and is a fundamental concept of quantum metrology [55, 56, 57, 58, 59]. The QFI plays an important role in quantum sensing and parameter estimation because it can provide a bound about the accuracy of quantum estimation, i.e., a larger QFI means a higher precision. Apart from quantum metrology, the QFI also connects to other aspects of quantum physics, such as quantum phase transition [60, 61, 62, 63], entanglement witness [64, 65, 66, 57, 67, 68], uncertainty relations [69], the calculation of quantum speedup limit time [70]. Another quantum version of Fisher information is related the notion of the WYSI information [71]. WYSI attracted a great deal of attention and is employed to understand entanglement [72, 73].
@@ II TIME DEPENDENT XY MODEL
To study the quench dynamics, we employ a linear ramp protocol in the absence of noise by setting the time‐dependent field to $h(t)=h_{0}(t)=t/\tau$, where $\tau$ is the ramp duration. To include noise, we superimpose a stochastic term, writing $h(t)=h_{0}(t)+R(t)=t/\tau+R(t)$, where $R(t)$ is a zero‐mean random fluctuation ($\langle R({t})\rangle=0$) confined to the interval $[t_{i},t_{f}]$. Therefore, the Hamiltonian of each mode can be decomposed into $H_{k}(t)=H_{0,k}(t)+R(t)H_{1,k}$ with $H_{1,k}=\sigma^{z}$. Here, we assume a white noise with Gaussian two-point correlations $\langle R(t)R(t^{\prime})\rangle=\xi^{2}\delta(t-t^{\prime})$, where $\xi$ characterizes the noise strength. Note that $\xi^{2}$ has units of time. White noise is approximately equivalent to the fast colored noise with exponentially decaying two-point correlations (Ornstein-Uhlenbeck process) [76]. Here, we assume the system initially in a large magnetic field background, namely $h_{i}{\ll}h_{c}{=}-1$, where the ground state of the system is paramagnetic. For this state all modes are initially in their lowest energy level, i.e., $|\psi(t=0)\rangle=\prod_{k>0}|\phi^{-}_{k}(t=0)\rangle$. Since the condition for an adiabatic dynamics breaks in the vicinity of the QCPs, after a ramp across the critical points $h_{c}=\pm 1$ to any final value $h(t)>h_{c}$, the final state $|\psi_{k}(t)\rangle$ does not remain in the groundstate of the final Hamiltonian. Therefore, the final state reads
@@
where $|\psi_{k}(t)\rangle=v_{k}|\phi^{-}_{k}(t)\rangle+u_{k}|\phi^{+}_{k}(t)\rangle$ with $|u_{k}|^{2}+|v_{k}|^{2}=1$, and $|u_{k}|^{2}$ is the probability of finding the $k$th mode in the excited state with energy $+\varepsilon_{k}$ at the end of the quench at time $t$. Hence any correlations among the system components that are obtained under such a ramp is a nonadiabatic quench process and must therefore include contributions from the excited states of the system. In the noiseless case, the amplitudes $u_{k}(t)$ and $v_{k}(t)$ are easily obtainable from the von Neumann equation
$\dot{\rho}_{0,k}(t)={d\over dt}\rho_{0,k}(t)=-i[H_{0,k}(t),\rho_{0,k}(t)].$
@@
In the presence of noise, the amplitudes $u_{k}(t)$ and $v_{k}(t)$ are obtained by numerically solving the exact master equation for the averaged density matrix $\rho_{k}(t)$ for a given mode $k$ and white-noise realization $R(t)$ [77, 78, 79, 76, 80, 81, 82, 83],
@@
where ${\cal H}_{0,k}(t)$ is the noise-free Hamiltonian while $R(t){\cal H}_{1}=R(t)\sigma^{z}$ expresses the “noisy” part for the full Hamiltonian ${\cal H}_{\xi,k}(t)={\cal H}_{0,k}(t)+R(t){\cal H}_{1}$. This master equation has the form of a von Neumann equation with an added term $-({\xi^{2}}/{2})[\sigma^{z},[\sigma^{z},\rho_{k}(t)]]$ representing the effect from the noise.
@@ III Quantum Fisher Information
Estimation theory is an important topic in different areas of physics [84, 85, 86]. In general phase estimation perspective, the evolution of a mixed quantum state, given by the density matrix $\varrho$, under a unitary transformation, can be described as $\varrho_{\theta}=e^{-iA\theta}\varrho e^{iA\theta}$, where $\theta$ is the phase shift and $A$ is an operator. The estimation accuracy for $\theta$ is bounded by the quantum Cramér-Rao inequality [87, 88]:
@@ IV Wigner-Yanase skew information
${\rm LQC}^{x}_{\ell,\ell+r}$ $=1-2(\alpha_{\varrho}\beta_{\varrho}+\gamma_{\varrho}\delta_{\varrho})-4{\rm Re}\Big{[}\lambda_{\varrho}\nu_{\varrho}\Big{]},$ (28)
${\rm LQC}^{y}_{\ell,\ell+r}$ $=1-2(\alpha_{\varrho}\beta_{\varrho}+\gamma_{\varrho}\delta_{\varrho})+4{\rm Re}\Big{[}\lambda_{\varrho}\nu_{\varrho}\Big{]},$
@@ V Numerical Results
Figure 1: Quantum Fisher information for nearest‐neighbour $F_{\ell,\ell+1}$ (a), and next‐nearest‐neighbour $F_{\ell,\ell+2}$ (b) as a function of the instantaneous field $h(t)$ during a linear ramp from $h_{i}=-30$ to $h_{f}$, in the absence of noise. Insets zoom into the critical region $-2<h(t)<2$. The lower panel show the Quantum Fisher information $F_{\ell,\ell+1}$ (c) and $F_{\ell,\ell+2}$ (d) versus ramp duration $\tau$ for a full quench from $h_{i}=-30$ to $h_{f}=30$.
@@
Figure 3: (a)-(c) Wigner–Yanase skew information between nearest‐neighbour spins ${\rm LQC}^{\alpha}_{\ell,\ell+1}$ for $\alpha=\{x,y,z\}$ as a function of the instantaneous field $h(t)$ during a noiseless ramp from $h_{i}=-30$ to $h(t)$. (d)-(f) Wigner–Yanase skew information for next‐nearest‐neighbour spins ${\rm LQC}^{\alpha}_{\ell,\ell+2}$ under the same ramp protocol and noiseless conditions. Insets zoom into the critical region $-2<h(t)<2$.
@@ V.1 Quantum Fisher Information
In our numerical investigation, we first explore the dynamics of the local quantum Fisher information under both noiseless and noisy linear quenches. Figures 1(a) and (b) show $F_{\ell,\ell+1}$ and $F_{\ell,\ell+2}$, respectively, as functions of the instantaneous field $h_{0}(t)$ for a chain of length $N=200$, with initial field $h_{i}=-30$ and several ramp durations $\tau$. As seen, despite to the entanglement, local quantum Fisher information is non-zero between nearest-neighbor spins. During the adiabatic evolution (very slow quench $\tau=500$), QFI closely follow the static values, i.e., those obtained from the ground state of the system at each value of $h_{0}(t)$, and assumes its minimum values at the critical points $h_{c}=\pm 1$, see Fig. 1(a) and (b). This behavior is expected from the adiabatic theorem [93, 94] and as a consequence of the finite size of the system. Indeed the energy gap closes as an inverse function of the system size, remaining nonzero for any finite size $L$, so that it is always possible to reach the adiabatic limit provided $\tau$ is long enough. More precisely, it has been shown that, the probability of having an adiabatic evolution at size $L$ is given by $p(\tau)=1-\exp(-2\pi^{3}\tau/L^{2})$ [46], so that the maximum quench rate at which the evolution is adiabatic decays as $1/L^{2}$. By decreasing $\tau$ the quench time to intermediate values ($2010<\tau<$ 20), QFI is nonzero at each value of $h_{0}(t)$, with small oscillations.
@@ V.2 Wigner-Yanas-Skew Information
On the other hand, the WYSI component ${\rm LQC}^{y}_{\ell,\ell+1}$ shows markedly different behaviour (Fig. 3(b)). Even in the static limit, it displays a clear minimum at $h_{0}(t)=0$ such as the fast quenches. While a pronounced maximum persists at $h_{0}(t)=1$, the feature at $h_{0}(t)=-1$ becomes only a kink rather than a true peak. Turning to Fig. 3(c), ${\rm LQC}^{z}_{\ell,\ell+1}$ mirrors the $x$–component’s approach to adiabaticity, with a pronounced maximum at $h_{0}(t)=0$ for slow quenches, the static limit, underscoring the dominant longitudinal correlations at zero field. As the quench rate increases, this central peak broadens and its amplitude diminishes, yet it persists as a maximum rather than inverting into a minimum. Furthermore, at high fields ($|h_{0}(t)|\gg 1$), ${\rm LQC}^{z}_{\ell,\ell+1}$ decays to zero, in stark contrast to the finite plateaus reached by the $x$ and $y$ components. Moreover, the next‐nearest‐neighbour component ${\rm LQC}^{\alpha}_{\ell,\ell+2}$ (Fig. 3(d)-(f)) closely mirrors the behaviour of its nearest‐neighbour counterpart. In Fig. 3(d)–(f) one again observes pronounced features at the critical fields: a central extremum at $h_{0}(t)=0$ that broadens and diminishes with increasing quench rate, inflections or secondary extrema near $h_{0}(t)=\pm 1$, and, for the $z$-component, a decay toward zero at large $|h_{0}(t)|$. Compared to the nearest‐neighbour curves, all peaks and plateaus are slightly reduced in amplitude, reflecting the weaker coherence between spins separated by two lattice sites.
To compare with our previous QFI results, we plot in Figs. 4(a)-(d), the Wigner–Yanase skew information versus ramp duration $\tau$ for a quench from $h_{i}=-30$ to $h_{f}=30$ for both nearest‐ and next‐nearest‐neighbour $x$ and $y$ components. We omit the $z$‐component here, as it is negligible at high fields and would add little to this analysis. In the noiseless scenario, the skew information increases monotonically with $\tau$, rising rapidly at small $\tau$ due to strong nonadiabatic excitations and then saturating to its static (adiabatic) value for large $\tau$. This behaviour holds for all four cases: nearest‐neighbour $x$ and $y$, and next‐nearest‐neighbour $x$ and $y$. In the presence of noise, however, the skew information no longer simply saturates. Instead, analogous to the QFI, each curve develops a maximum at an intermediate $\tau_{m}(\xi)$ and then decreases for $\tau>\tau_{m}(\xi)$. Thus, all four $x$ and $y$ components, whether nearest‐ or next‐nearest neighbours, display qualitatively the same approach to QFI, with the next‐nearest‐neighbour curves slightly attenuated in amplitude faster by increasing $\tau$, for higher noise intensity.

### v3

@@ Abstract
We study the effect of noise on the dynamics of the transverse-field Ising model quenched across a quantum critical point. To quantify two-spin correlations, we employ the quantum Fisher information (QFI) and the Wigner–Yanase skew information (WYSI) as measures of quantum coherence. In the noiseless case, in contrast to the dynamics of entanglement in anisotropic XY chains, both QFI and WYSI increase monotonically with the ramp quench time, approaching their adiabatic limits without exhibiting any Kibble–Zurek–type scaling with quench duration. In contrast, when noise is added to the quench protocol, the coherence dynamics change qualitatively: QFI and WYSI both decay exponentially with the time scale of a ramp quench, with an exponent determined by the noise intensity. Furthermore, the maximum ramp time, at which either of these measures reach their maximum, scales linearly with the noise variance, featuring the same exponent that determines the optimal annealing time for minimizing defect production in noisy quantum annealing.
@@ I Introduction
Quantum phase transitions (QPTs) have been one of the most important topics in the area of strongly correlated systems [1, 2, 3]. There have been recent developments in theoretical investigation of QPTs from the viewpoint of quantum information theory. Quantities like concurrence [4, 5], entanglement entropy [6, 7], fidelity susceptibility [8, 9, 10, 11], and geometric phases [12, 13] have been shown to capture the non-analyticities associated with a quantum critical point (QCP). Recent advances in the studies of ultracold atoms trapped in optical lattices have opened a new avenue of investigation of the non-equilibrium dynamics of QPTs in quantum many body systems [14, 15]. Divergences in correlation length and time scales at QCPs introduce non-trivial effects in the dynamics of a quantum system near a QCP. In particular, if a parameter of the Hamiltonian is tuned slowly so as to drive the system through a QCP, the dynamics fails to be adiabatic for any finite rate of variation of the parameter. This non-adiabaticity in the response of the system eventually leads to generation of defects in the final state. A quantitative analysis of defect production in a system following a quench can be carried out using the Kibble-Zurek (KZ) theory [16, 17] extended to quantum spin chains [18, 19, 20, 21, 22, 23, 24, 25, 26].
@@
This raises a fundamental question about the generation of quantum correlations after a noisy quench across a QCP. If such correlations are generated, how does noise influence their dynamical behavior? To address this, we analyze the dynamics of the quantum Fisher information (QFI) and the Wigner–Yanase skew information (WYSI) in the transverse-field Ising model (TFI) under both noiseless and noisy driving. QFI [56] is an extension of Fisher information [57] to the quantum realm, and is a fundamental concept of quantum metrology [58, 59, 60, 61, 62]. The QFI plays an important role in quantum sensing and parameter estimation because it can provide a bound about the accuracy of quantum estimation, i.e., a larger QFI means a higher precision. Apart from quantum metrology, the QFI also connects to other aspects of quantum physics, such as quantum phase transition [63, 64, 65, 66], entanglement witness [67, 68, 69, 60, 70, 71], uncertainty relations [72], the calculation of quantum speedup limit time [73]. Another quantum version of Fisher information is the WYSI [74], which has attracted considerable attention and is used to investigate entanglement [75, 76].
@@ II TIME DEPENDENT XY MODEL
To study the quench dynamics, we employ a linear ramp protocol in the absence of noise by setting the time‐dependent field to $h(t)=h_{0}(t)=t/\tau$, where $\tau$ is the ramp duration. To include noise, we superimpose a stochastic term, writing $h(t)=h_{0}(t)+R(t)=t/\tau+R(t)$, where $R(t)$ is a zero‐mean random fluctuation ($\langle R({t})\rangle=0$) confined to the interval $t_{i},t_{f}]$. Therefore, the Hamiltonian of each mode can be decomposed into $H_{k}(t)=H_{0,k}(t)+R(t)H_{1,k}$ with $H_{1,k}=\sigma^{z}$. Here, we assume a white noise with Gaussian two-point correlations $\langle R(t)R(t^{\prime})\rangle=\xi^{2}\delta(t-t^{\prime})$, where $\xi$ characterizes the noise strength. Note that $\xi^{2}$ has units of time. White noise is approximately equivalent to the fast colored noise with exponentially decaying two-point correlations (Ornstein-Uhlenbeck process) [79]. Here, we assume the system initially in a large magnetic field background, namely $h_{i}\ll h_{c}=-1$, where the ground state of the system is paramagnetic. For this state all modes are initially in their lowest energy level, i.e., $|\psi(t=0)\rangle=\prod_{k>0}|\phi^{-}_{k}(t=0)\rangle$. Since the condition for an adiabatic dynamics breaks in the vicinity of the QCPs, after a ramp across the critical points $h_{c}=\pm 1$ to any final value $h(t)>h_{c}$, the final state $|\psi_{k}(t)\rangle$ does not remain in the groundstate of the final Hamiltonian. Therefore, for the noiseless case the final state reads
@@
where $|\psi_{k}(t)\rangle=v_{k}|\phi^{-}_{k}(t)\rangle+u_{k}|\phi^{+}_{k}(t)\rangle$ with $|u_{k}|^{2}+|v_{k}|^{2}=1$, and $|u_{k}|^{2}$ is the probability of finding the $k$th mode in the excited state with energy $\varepsilon_{k}$ at the end of the quench at time $t$. Hence any correlations among the system components that are obtained under such a ramp is a nonadiabatic quench process and must therefore include contributions from the excited states of the system. In the noiseless case, the amplitudes $u_{k}(t)$ and $v_{k}(t)$ are easily obtainable from the von Neumann equation
$\dot{\rho}_{0,k}(t)={d\over dt}\rho_{0,k}(t)=-i[H_{0,k}(t),\rho_{0,k}(t)].$ (6)
@@
In the presence of noise, the averaged density matrix $\rho_{k}(t)$ for a given mode $k$ and white-noise realization $R(t)$ are obtained by numerically solving the exact master equation for [80, 81, 82, 79, 83, 84, 85],
@@
where ${\cal H}_{0,k}(t)$ is the noise-free Hamiltonian while $R(t){\cal H}_{1}=R(t)\sigma^{z}$ expresses the “noisy” part for the full Hamiltonian ${\cal H}_{\xi,k}(t)={\cal H}_{0,k}(t)+R(t){\cal H}_{1}$. This master equation has the form of a von Neumann equation with an added term $({\xi^{2}}/{2})[\sigma^{z},[\sigma^{z},\rho_{k}(t)]]$ representing the effect from the noise.
@@ III Quantum Fisher Information
Estimation theory is an important topic in different areas of physics [86, 87, 88, 89, 90]. In general phase estimation perspective, the evolution of a mixed quantum state, given by the density matrix $\varrho$, under a unitary transformation, can be described as $\varrho_{\theta}=e^{-iA\theta}\varrho e^{iA\theta}$, where $\theta$ is the phase shift and $A$ is an operator. The estimation accuracy for $\theta$ is bounded by the quantum Cramér-Rao inequality [91, 92]:
@@ IV Wigner-Yanase skew information
${\rm LQC}^{x}_{\ell,\ell+r}$ $=1-2(\alpha_{\varrho}\beta_{\varrho}+\gamma_{\varrho}\delta_{\varrho})-4{\mathrm{R}e}\Big[\lambda_{\varrho}\nu_{\varrho}\Big],$ (29)
${\rm LQC}^{y}_{\ell,\ell+r}$ $=1-2(\alpha_{\varrho}\beta_{\varrho}+\gamma_{\varrho}\delta_{\varrho})+4{\mathrm{R}e}\Big[\lambda_{\varrho}\nu_{\varrho}\Big],$
@@ V Numerical Results
Figure 1: Quantum Fisher information for nearest‐neighbour $F_{\ell,\ell+1}$ (a), and next‐nearest‐neighbour $F_{\ell,\ell+2}$ (b) as a function of the instantaneous field $h(t)$ during a linear ramp from $h_{i}=-30$ to $h_{f}$, in the absence of noise. Insets zoom into the critical region 2<h(t)<2-2<h(t)<2. The lower panel show the Quantum Fisher information $F_{\ell,\ell+1}$ (c) and $F_{\ell,\ell+2}$ (d) versus ramp duration $\tau$ for a full quench from $h_{i}=-30$ to $h_{f}=30$.
@@
Figure 3: (a)-(c) Wigner–Yanase skew information between nearest‐neighbour spins ${\rm LQC}^{\alpha}_{\ell,\ell+1}$ for $\alpha=\{x,y,z\}$ as a function of the instantaneous field $h(t)$ during a noiseless ramp from $h_{i}=-30$ to $h(t)$. (d)-(f) Wigner–Yanase skew information for next‐nearest‐neighbour spins ${\rm LQC}^{\alpha}_{\ell,\ell+2}$ under the same ramp protocol and noiseless conditions. Insets zoom into the critical region 2<h(t)<2-2<h(t)<2.
@@ V.1 Quantum Fisher Information
In our numerical investigation, we first explore the dynamics of the local quantum Fisher information under both noiseless and noisy linear quenches. Figures 1(a) and (b) show $F_{\ell,\ell+1}$ and $F_{\ell,\ell+2}$, respectively, as functions of the instantaneous field $h_{0}(t)$ for a chain of length $N=200$, with initial field $h_{i}=-30$ and several ramp durations $\tau$. In contrast to the dynamics of entanglement in anisotropic XY chains, local quantum Fisher information is non-zero between nearest-neighbor spins. During the adiabatic evolution (very slow quench $\tau=500$), QFI closely follow the static values, i.e., those obtained from the ground state of the system at each value of $h_{0}(t)$, and assumes its minimum values at the critical points $h_{c}=\pm 1$, see Fig. 1(a) and (b). This behavior is expected from the adiabatic theorem [97, 98] and as a consequence of the finite size of the system. Indeed the energy gap closes as an inverse function of the system size, remaining nonzero for any finite size $L$, so that it is always possible to reach the adiabatic limit provided $\tau$ is long enough. More precisely, it has been shown that, the probability of having an adiabatic evolution at size $L$ is given by $p(\tau)=1-\exp(-2\pi^{3}\tau/L^{2})$ [49], so that the maximum quench rate at which the evolution is adiabatic decays as $1/L^{2}$. By decreasing $\tau$ the quench time to intermediate values ($10<\tau<20$), QFI is nonzero at each value of $h_{0}(t)$, with small oscillations.
@@ V.2 Wigner-Yanas-Skew Information
On the other hand, the WYSI component ${\rm LQC}^{y}_{\ell,\ell+1}$ shows markedly different behaviour (Fig. 3(b)). Even in the static limit, it displays a clear minimum at $h_{0}(t)=0$ such as the fast quenches. While a pronounced maximum persists at $h_{0}(t)=1$, the feature at $h_{0}(t)=-1$ becomes only a kink rather than a true peak. Turning to Fig. 3(c), ${\rm LQC}^{z}_{\ell,\ell+1}$ mirrors the xx–component’s approach to adiabaticity, with a pronounced maximum at $h_{0}(t)=0$ for slow quenches, the static limit, underscoring the dominant longitudinal correlations at zero field. As the quench rate increases, this central peak broadens and its amplitude diminishes, yet it persists as a maximum rather than inverting into a minimum. Furthermore, at high fields ($|h_{0}(t)|\gg 1$), ${\rm LQC}^{z}_{\ell,\ell+1}$ decays to zero, in stark contrast to the finite plateaus reached by the $x$ and $y$ components. Moreover, the next‐nearest‐neighbour component ${\rm LQC}^{\alpha}_{\ell,\ell+2}$ (Fig. 3(d)-(f)) closely mirrors the behaviour of its nearest‐neighbour counterpart. In Fig. 3(d)–(f) one again observes pronounced features at the critical fields: a central extremum at $h_{0}(t)=0$ that broadens and diminishes with increasing quench rate, inflections or secondary extrema near $h_{0}(t)=\pm 1$, and, for the $z$-component, a decay toward zero at large $|h_{0}(t)|$. Compared to the nearest‐neighbour curves, all peaks and plateaus are slightly reduced in amplitude, reflecting the weaker coherence between spins separated by two lattice sites.
To compare with our previous QFI results, we plot in Figs. 4(a)-(d), the Wigner–Yanase skew information versus ramp duration $\tau$ for a quench from $h_{i}=-30$ to $h_{f}=30$ for both nearest‐ and next‐nearest‐neighbour $x$ and $y$ components. We omit the zz‐component here, as it is negligible at high fields and would add little to this analysis. In the noiseless scenario, the skew information increases monotonically with $\tau$, rising rapidly at small $\tau$ due to strong nonadiabatic excitations and then saturating to its static (adiabatic) value for large $\tau$. This behaviour holds for all four cases: nearest‐neighbour $x$ and $y$, and next‐nearest‐neighbour $x$ and $y$. In the presence of noise, however, the skew information no longer simply saturates. Instead, analogous to the QFI, each curve develops a maximum at an intermediate $\tau_{m}(\xi)$ and then decreases for $\tau>\tau_{m}(\xi)$. Thus, all four $x$ and $y$ components, whether nearest‐ or next‐nearest neighbours, display qualitatively the same approach to QFI, with the next‐nearest‐neighbour curves slightly attenuated in amplitude faster by increasing $\tau$, for higher noise intensity.